# The radius of statistical efficiency

### v1

@@ A.2 Spectral sets and functions of measures
$=\left\|\sqrt{\lambda(A)}-\sqrt{\lambda(B)}\right\|^{2}_{2},$ (56)
where (50) follows from Lemma A.9, the estimate (51) follows from expanding the Frobenius norm, (52) uses the variable substitution $Z=V^{\top}U$, the estimate (53) follows from von Neumann’s trace inequality (Lemma A.8), and (54) uses the fact that eigenvalues and singular values coincide for PSD matrices. Taking the infimum over $V\in O(d)$ shows the claimed inequality $\geq$ in (48).
Next, the fact that any matrix in (49) is a nearest point of $\mathcal{O}(B)$ to $A$ follows directly from the expression (48). To see the converse, observe that $VBV\in O(B)$ is the closest point to $A$ if and only if the chain of inequalities (50)-(56) holds as equalities. Since the only inequality appears in (53), applying Lemma A.8 we see that equality holds if and only if there exist matrices $M_{1},M_{2},Z\in O(d)$ such that
@@
In particular, multiplying each equation by its transpose yields the expressions
@@
thereby concluding the proof. ∎
We are now ready to complete the proof of Theorem A.2 in the Gaussian setting. In the proof, we will use the basic fact that for any vectors $v,w\in\mathbf{R}^{d}$, the inequality holds:
@@
Theorem A.12 (Envelope and prox-map in Gaussian space).
Consider a symmetric function $f\colon\mathbf{R}^{d}_{+}\to\mathbf{R}\cup\{+\infty\}$ and define the function on PSD matrices $F\colon{\mathcal{S}}^{d}_{+}\to\mathbf{R}\cup\{+\infty\}$ by setting $F(A)=f\left(\lambda\left(A\right)\right).$ Then, equality holds:
$F_{\rho}(A)=f_{\rho}(\lambda(A))\qquad\forall A\in\bm{S}^{d}_{+}.$
Moreover, the following expression holds:
$\operatorname{prox}_{\rho F}(A)=\{U\operatorname{Diag}(v)U^{\top}:v\in\operatorname{prox}_{\rho f}(\lambda(A)),A=U\operatorname{Diag}(\lambda(A))U^{\top},~{}U\in O(d)\}.$
@@
For any matrix $A\in\bm{S}^{d}_{+}$, we successively compute
$F_{\rho}(A)$ $=\inf_{B\succeq 0}~{}F(A)+\tfrac{1}{2\rho}W^{2}_{2}(A,B)$ (59)
@@
$=\inf_{v\geq 0}~{}f(v)+\tfrac{1}{2\rho}\|\sqrt{\lambda(A)}-\sqrt{v^{\uparrow}}\|^{2}_{2}$ (63)
@@
where (63) follows from Lemma A.11 and (64) follows from (58). Next, observe from the chain of equalities that $B=U\operatorname{Diag}(v)U^{\top}$ lies in $\operatorname{prox}_{\rho F}(A)$ for some $U\in O(d)$ if, and only if, $v$ lies in $\operatorname{prox}_{\rho f}(\lambda(A))$ and equality $d_{2}(A,B)=d_{2}(A,\mathcal{O}(\operatorname{Diag}(v)))$ holds. Appealing to Lemma A.11, this equality holds if and only if we may write $A=U\operatorname{Diag}(\lambda(A))U^{\top}$. Thus the proof is complete. ∎
We now move on to establishing Theorem A.2 in full generality, i.e. outside the Gaussian setting. We begin by extending Lemma A.11.
Lemma A.13 (Distance to orbit: general case).
Fix a matrix $B\in\bm{S}^{d}_{+}$ and define the set of measures
:=$\mathcal{M}:=\left\{\nu\in\mathcal{P}_{2}:~{}\Sigma_{\nu}\in\mathcal{O}(B)\right\}.$
Then any zero-mean measure $\mu\in\mathcal{P}_{2}$ satisfies:
$W_{2}^{2}(\mu,\mathcal{M})=W_{2}^{2}(\Sigma_{\mu},\mathcal{O}(B)).$ (66)
@@
We suppose first that $\Sigma_{\mu}$ is positive definite. Observe now
$W_{2}(\mu,\mathcal{M})$ $=\inf_{\pi\in\Pi(\mu,\nu),\,\nu\in\mathcal{M}}\operatorname{\mathbb{E}}_{\pi}\|x-y\|^{2}_{2}$
$\geq\inf_{\operatorname{\mathbb{E}}[xx^{\top}]=\Sigma_{\mu},~{}\operatorname{\mathbb{E}}[yy^{\top}]\in\mathcal{O}(B)}\operatorname{\mathbb{E}}\|x-y\|^{2}_{2}=d_{2}(\Sigma_{\mu},\mathcal{O}(B)),$
where the inequality follows from Lemma A.10. To see the reverse inequality, consider an eigenvalue decomposition $\Sigma_{\mu}=V\operatorname{Diag}(\lambda(\Sigma_{\mu}))V^{\top}$ for some $V\in O(d)$. Define now the matrix $\bar{B}:=V{\rm Diag}(\lambda(B))V^{\top}$ and set $T:=\bar{B}^{1/2}\Sigma_{\mu}^{-1/2}$. Then, clearly the measure $\nu:=T_{\#}\mu$ satisfies $\operatorname{\mathbb{E}}_{z\sim\nu}[zz^{\top}]=T\operatorname{\mathbb{E}}_{\mu}[xx^{\top}]T^{\top}=T\Sigma_{\mu}T^{\top}=\bar{B}$ and therefore $T_{\#}\mu$ lies in $\mathcal{M}$. Thus, we conclude
$W_{2}^{2}(\mu,\mathcal{M})\leq W_{2}^{2}(\mu,\nu)$ $=\operatorname{\mathbb{E}}_{\mu}\|x-Tx\|^{2}_{2}$
$=\operatorname{\mathbb{E}}_{\mu}\|x\|^{2}-2\operatorname{\mathbb{E}}_{\mu}\operatorname{tr}(Txx^{\top})+\operatorname{\mathbb{E}}_{\mu}\operatorname{tr}(T^{\top}Txx^{\top})$
$=\operatorname{tr}(\Sigma_{\mu})-2\operatorname{tr}(\bar{B}^{1/2}\Sigma_{\mu}^{1/2})+\operatorname{tr}(\bar{B})$
$=\sum_{i=1}^{d}\left(\sqrt{\lambda_{i}(\Sigma_{\mu})}-\sqrt{\lambda_{i}(B)}\right)^{2},$
Using Lemma A.11 yields the claimed expression (66).
Finally, if $\Sigma_{\mu}$ is not positive definite, then $\mu$ is almost surely supported on some subspace $\mathcal{L}$. We now define the perturbed distribution $\mu_{t}=\mu\times g_{t}$ where $g_{t}$ is a zero-mean Gaussian distribution supported on $\mathcal{L}^{\perp}$ with covariance $t\cdot I$. Then, by Lemma A.2, $W_{2}(\mu_{t},\mu)=W_{2}(\mu_{t},(\mathsf{P}_{\mathcal{L}})_{\#}\mu_{t})\rightarrow 0$, which in turn implies $\operatorname{\mathbb{E}}_{{\mu_{t}}}xx^{\top}\rightarrow\operatorname{\mathbb{E}}_{\mu}xx^{\top}.$ From (66) we have $W_{2}^{2}(\mu_{t},\mathcal{M})=W_{2}^{2}(\operatorname{\mathbb{E}}_{\mu_{t}}xx^{\top},\mathcal{O}(B))$. Letting $t$ go to zero, we deduce the desired equality $W_{2}^{2}(\mu,\mathcal{M})=W_{2}^{2}(\Sigma_{\mu},\mathcal{O}(B))$. ∎
The proof of Theorem A.2 now proceeds in exactly the same way as that of Theorem A.12, with Lemma A.13 being used instead of Lemma A.11.
@@
The argument is essentially the same as in Theorem A.12. We detail it here for completeness. Define the matrix $A:=\Sigma_{\mu}$. We successively compute
$F_{\rho}(\mu)$ $=\inf_{\nu\in\mathcal{P}_{2}}~{}F(\nu)+\tfrac{1}{2\rho}W^{2}_{2}(\mu,\nu)$ (67)
$=\inf_{u\geq 0}\inf_{\nu:\,\Sigma_{\nu}\in\mathcal{O}(\operatorname{Diag}(u))}~{}f(u)+\tfrac{1}{2\rho}W^{2}_{2}(\mu,\nu)$ (68)
$=\inf_{u\geq 0}~{}f(u)+\tfrac{1}{2\rho}\inf_{\nu:\,\Sigma_{\nu}\in\mathcal{O}(\operatorname{Diag}(u))}W^{2}_{2}(\mu,\nu)$ (69)
$=\inf_{u\geq 0}~{}f(u)+\tfrac{1}{2\rho}W_{2}^{2}(A,\mathcal{O}(\operatorname{Diag}(u)))$ (70)
$=\inf_{v\geq 0}~{}f(v)+\tfrac{1}{2\rho}\|\sqrt{\lambda(A)}-\sqrt{u^{\uparrow}}\|^{2}_{2}$ (71)
@@
where (70) follows from Lemma A.13, the estimate (71) follows from Lemma A.11, and (72) follows from (58). This completes the proof. ∎
@@
The results presented in this section have close analogues in the existing literature in matrix analysis and optimization. Namely a function $F\colon\bm{S}^{d}\to\mathbf{R}\cup\{+\infty\}$ is called orthogonally invariant (or spectral) if the equality holds:
@@
Evidently such functions are fully described by their restriction to diagonal matrices. More precisely, a function $F$ is orthogonally invariant if and only if there exists a symmetric function $f\colon\mathbf{R}^{d}\to\mathbf{R}\cup\{+\infty\}$ satisfying $F(X)=f(\lambda(X))$. A pervasive theme in the study of such functions is that various variational properties of the permutation-invariant function $f$ are inherited by the induced spectral function $F=f\circ\lambda$; see e.g. [42, 72, 40, 41, 23, 22, 60, 25]. For example, $f$ convex if and only if $F$ is convex [25, 40], $f$ is $C^{p}$-smooth if and only of $F$ is $C^{p}$-smooth [60, 72, 63], and so forth. A useful result in this area is the expression for the Moreau envelope obtained in [22, 33]
@@
For example, as explained in [33], it readily yields expressions relating generalized derivatives of $f$ and $F$. Crucially, in (74) the Moreau envelope $F_{\rho}$ is computed with respect to the Frobenius norm on $\bm{S}^{d}$ and the Moreau envelope $f_{\rho}$ is computed with respect to the $\ell_{2}$ norm on $\mathbf{R}^{d}$. Thus the results presented in the section extend this circle of ideas to functions defined on the Wasserstein space.
@@ Appendix B Proofs from Section 2
Suppose otherwise that there exists a sequence $\nu_{i}\in\mathcal{Q}^{\prime}$ with $W_{2}(\nu_{i},\mu)\leq r$ satisfying ${\rm REG}(\nu_{i})\to\infty$. Then from (11) we deduce $W_{2}(\nu_{i},\mathcal{E})={\rm RSE}(\nu_{i})\to 0$. Subsequently, using the triangle inequality yields ${\rm RSE}(\mu)=W_{2}(\mu,\mathcal{E})\leq W_{2}(\nu_{i},\mu)+W_{2}(\nu_{i},\mathcal{E})$. Letting $i$ tend to infinity, we deduce $W_{2}(\mu,\mathcal{E})\leq r$, which is a contradiction. Thus no such sequence $\nu_{i}$ exists and (12) holds for some $M>0$.
Suppose now that for some $q>0$, the inequality $c\geq{\rm REG}(\nu)\cdot{\rm RSE}(\nu)^{q}$ holds. Then, similarly as above, the triangle inequality for any $\nu\in\mathcal{Q}^{\prime}\setminus\mathcal{E}$ yields
@@
Rearranging yields the estimate ${\rm REG}(\nu)\leq c({\rm RSE}(\mu)-r)^{-1/q}$, thereby completing the proof.

### v2

@@ A.2 Spectral sets and functions of measures
$=\Big\|\sqrt{\lambda(A)}-\sqrt{\lambda(B)}\Big\|^{2}_{2},$ (59)
where (53) follows from Lemma A.9, the equality (54) follows from expanding the Frobenius norm, (55) uses the variable substitution $Z=V^{\top}U$, the inequality (56) follows from von Neumann’s trace inequality (Lemma A.8), and (57) uses the fact that eigenvalues and singular values coincide for PSD matrices. Taking the infimum over $V\in O(d)$ shows the claimed inequality $\geq$ in (51).
Next, the fact that any matrix in the set (52) is a nearest point of $\mathcal{O}(B)$ to $A$ follows directly from the expression (51). To see the converse, let $V\in O(d)$ and observe from (53)–(59) that $VBV^{\top}$ is a nearest point of $\mathcal{O}(B)$ to $A$ if and only if (56) holds with equality; applying Lemma A.8, we see that equality holds if and only if there exist matrices $M_{1},M_{2},Z\in O(d)$ such that
@@
In particular, right-multiplying each equation by its transpose yields the expressions
@@
thereby completing the proof. ∎
We are now ready to complete the proof of Theorem A.2 in the Gaussian setting. In the proof, we will use the basic fact that for any vectors $v,w\in\mathbf{R}^{d}$, we have
@@
Theorem A.12 (Envelope and prox-map on Gaussian space).
Consider a symmetric function $f\colon\mathbf{R}^{d}_{+}\to\mathbf{R}\cup\{+\infty\}$ and define the function on PSD matrices $F\colon\bm{S}^{d}_{+}\to\mathbf{R}\cup\{+\infty\}$ by setting $F(A)=f(\lambda(A)).$ Then, for any $\rho>0$ and $A\in\bm{S}^{d}_{+}$, the following equalities hold:
$F_{\rho}(A)=f_{\rho}(\lambda(A))$
and
$\operatorname{prox}_{\rho F}(A)=\{U\operatorname{Diag}(v)U^{\top}\mid v\in\operatorname{prox}_{\rho f}(\lambda(A)),~A=U\operatorname{Diag}(\lambda(A))U^{\top},~U\in O(d)\}.$
@@
Given $\rho>0$ and $A\in\bm{S}^{d}_{+}$, we successively compute
$F_{\rho}(A)$ $=\inf_{B\succeq 0}~F(B)+\tfrac{1}{2\rho}W^{2}_{2}(A,B)$ (62)
@@
$=\inf_{v\geq 0}~f(v)+\tfrac{1}{2\rho}\Big\|\sqrt{\lambda(A)}-\sqrt{v^{\uparrow}}\Big\|^{2}_{2}$ (66)
@@
where (66) follows from Lemma A.11 and (67) follows from (61) and the symmetry of $f$. Next, observe from the chain of equalities above that $B=U\operatorname{Diag}(v)U^{\top}$ lies in $\operatorname{prox}_{\rho F}(A)$ for some $U\in O(d)$ if, and only if, $v$ lies in $\operatorname{prox}_{\rho f}(\lambda(A))$ and the equality $W_{2}(A,B)=W_{2}(A,\mathcal{O}(\operatorname{Diag}(v)))$ holds; appealing to Lemma A.11, this last equality holds if and only if we may write $A=U\operatorname{Diag}(\lambda(A))U^{\top}$. Thus the proof is complete. ∎
We now move on to establishing Theorem A.2 in full generality, i.e., outside the Gaussian setting. To begin, we establish an analogue of Lemma A.11.
Lemma A.13.
Let $\mathcal{C}\subset\bm{S}^{d}_{+}$ be a set of positive semidefinite matrices, and consider the set $\mathcal{P}_{\mathcal{C}}:=\{\nu\in\mathcal{P}_{2}^{\circ}\mid\Sigma_{\nu}\in\mathcal{C}\}.$ Then for any $\mu\in\mathcal{P}^{\circ}_{2}$, the following equality holds:
$W_{2}(\mu,\mathcal{P}_{\mathcal{C}})=W_{2}(\Sigma_{\mu},\mathcal{C}).$ (69)
@@
Note first that for any $\mu\in\mathcal{P}_{2}$, we have
$W^{2}_{2}(\mu,\mathcal{P}_{\mathcal{C}})=\inf_{\pi\in\Pi(\mu,\nu),\,\nu\in\mathcal{P}_{\mathcal{C}}}\mathop{\operatorname{\mathbb{E}}}_{(x,y)\sim\pi}\|x-y\|^{2}_{2}$ $\geq\inf_{x,y:\,\operatorname{\mathbb{E}}[xx^{\top}]=\Sigma_{\mu},\,\operatorname{\mathbb{E}}[yy^{\top}]\in\mathcal{C}}\operatorname{\mathbb{E}}\|x-y\|^{2}_{2}$
$=W_{2}^{2}(\Sigma_{\mu},\mathcal{C}),$
where the last equality follows from Lemma A.10. This establishes the inequality $\geq$ in (69). To see the reverse inequality, let $\mu\in\mathcal{P}^{\circ}_{2}$ and $\Sigma\in\mathcal{C}$. Suppose first that $\Sigma_{\mu}$ is invertible. In this case, the optimal transport map from $\mathsf{N}(0,\Sigma_{\mu})$ to $\mathsf{N}(0,\Sigma)$ is given by a linear function $T$, as in [51, Section 1.6.3]. Consider the pushforward measure $\bar{\nu}:=T_{\#}\mu$. Since the integral of a quadratic function depends on a probability measure only through its first two moments, it follows immediately from the linearity of $T$ that the first two moments of $T_{\#}\mu=\bar{\nu}$ and $T_{\#}\mathsf{N}(0,\Sigma_{\mu})=\mathsf{N}(0,\Sigma)$ coincide, so $\bar{\nu}\in\mathcal{P}_{\mathcal{C}}$. Likewise, ${\operatorname{\mathbb{E}}}_{\mu}\|x-Tx\|^{2}_{2}={\operatorname{\mathbb{E}}}_{\mathsf{N}(0,\Sigma_{\mu})}\|x-Tx\|^{2}_{2}$, so we conclude
$W^{2}_{2}(\mu,\mathcal{P}_{\mathcal{C}})\leq W_{2}^{2}(\mu,\bar{\nu})\leq{\operatorname{\mathbb{E}}}_{\mu}\|x-Tx\|^{2}_{2}=W_{2}^{2}(\mathsf{N}(0,\Sigma_{\mu}),\mathsf{N}(0,\Sigma))=W_{2}^{2}(\Sigma_{\mu},\Sigma).$
Taking the infimum over $\Sigma\in\mathcal{C}$ yields $W^{2}_{2}(\mu,\mathcal{P}_{\mathcal{C}})\leq W_{2}^{2}(\Sigma_{\mu},\mathcal{C})$, as desired.
Finally, if $\Sigma_{\mu}$ is not invertible, then $\mu$ is supported on the proper subspace $\mathcal{H}:=\operatorname{range}(\Sigma_{\mu})$ of $\mathbf{R}^{d}$. We then define the perturbed distribution $\mu_{t}:=\mu|_{\mathcal{H}}\times\gamma_{t}$ on $\mathbf{R}^{d}$, where $\gamma_{t}$ is the zero-mean Gaussian distribution supported on $\mathcal{H}^{\perp}=\ker(\Sigma_{\mu})$ with covariance $t\cdot I$. Now, Lemma A.2 shows $W_{2}^{2}(\mu_{t},\mu)=W_{2}^{2}(\mu_{t},(\mathsf{P}_{\mathcal{H}})_{\#}\mu_{t})=\operatorname{tr}\Sigma_{\gamma_{t}}\rightarrow 0$ as $t\rightarrow 0$, and clearly $\Sigma_{\mu_{t}}$ is invertible with $\Sigma_{\mu_{t}}\rightarrow\Sigma_{\mu}$ as $t\rightarrow 0$. From (69), we have $W_{2}(\mu_{t},\mathcal{P}_{\mathcal{C}})=W_{2}(\Sigma_{\mu_{t}},\mathcal{C})$. Letting $t\rightarrow 0$, we deduce $W_{2}(\mu,\mathcal{P}_{\mathcal{C}})=W_{2}(\Sigma_{\mu},\mathcal{C})$ by continuity. ∎
The proof of Theorem A.2 now proceeds in exactly the same way as that of Theorem A.12, with Lemma A.13 being used in conjunction with Lemma A.11.
@@
For any set $\mathcal{C}\subset\bm{S}^{d}_{+}$, define $\mathcal{P}_{\mathcal{C}}:=\{\nu\in\mathcal{P}_{2}^{\circ}\mid\Sigma_{\nu}\in\mathcal{C}\}$. Then for any $\rho>0$ and $\mu\in\mathcal{P}^{\circ}_{2}$, we successively compute
$F_{\rho}(\mu)$ $=\inf_{\nu\in\mathcal{P}_{2}^{\circ}}~F(\nu)+\tfrac{1}{2\rho}W^{2}_{2}(\mu,\nu)$ (70)
$=\inf_{u\geq 0}\,\inf_{\nu\in\mathcal{P}_{\mathcal{O}(\operatorname{Diag}(u))}}~f(u)+\tfrac{1}{2\rho}W^{2}_{2}(\mu,\nu)$ (71)
$=\inf_{u\geq 0}~f(u)+\tfrac{1}{2\rho}+W^{2}_{2}\big(\mu,\mathcal{P}_{\mathcal{O}(\operatorname{Diag}(u))}\big)$ (72)
$=\inf_{u\geq 0}~f(u)+\tfrac{1}{2\rho}W_{2}^{2}(\Sigma_{\mu},\mathcal{O}(\operatorname{Diag}(u)))$ (73)
$=\inf_{u\geq 0}~f(u)+\tfrac{1}{2\rho}\Big\|\sqrt{\lambda(\Sigma_{\mu})}-\sqrt{u^{\uparrow}}\Big\|^{2}_{2}$ (74)
@@
where (73) follows from Lemma A.13, the equality (74) follows from Lemma A.11, and (75) follows from (61) and the symmetry of $f$. This completes the proof. ∎
@@
The results presented in this section have close analogues in the existing literature on matrix analysis and optimization. Namely, a function $F\colon\bm{S}^{d}\to\mathbf{R}\cup\{+\infty\}$ is called orthogonally invariant (or spectral) if the equality holds:
@@
Evidently, such functions are fully described by their restriction to diagonal matrices. More precisely, a function $F$ is orthogonally invariant if and only if there exists a symmetric function $f\colon\mathbf{R}^{d}\to\mathbf{R}\cup\{+\infty\}$ satisfying $F(X)=f(\lambda(X))$. A pervasive theme in the study of such functions is that various variational properties of the permutation-invariant function $f$ are inherited by the induced spectral function $F=f\circ\lambda$; see, e.g., [41, 65, 43, 44, 22, 23, 63, 25]. For example, $f$ is convex if and only if $F$ is convex [25, 43], $f$ is $C^{p}$-smooth if and only of $F$ is $C^{p}$-smooth [63, 65, 67], and so forth. A useful result in this area is the expression for the Moreau envelope obtained in [23, 34]:
@@
For example, as explained in [34], it readily yields expressions relating generalized derivatives of $f$ and $F$. Crucially, in (77) the Moreau envelope $F_{\rho}$ is computed with respect to the Frobenius norm on $\bm{S}^{d}$ and the Moreau envelope $f_{\rho}$ is computed with respect to the $\ell_{2}$ norm on $\mathbf{R}^{d}$. Thus the results presented in the section extend this circle of ideas to functions defined on the Wasserstein-2 space.
@@ Appendix B Proofs from Section 2
Suppose otherwise that there exists a sequence $\nu_{i}\in\mathcal{Q}^{\prime}$ with $W_{2}(\nu_{i},\mu)\leq r$ satisfying ${\rm REG}(\nu_{i})\to\infty$. Then from (11) we deduce $W_{2}(\nu_{i},\mathcal{E})={\rm RSE}(\nu_{i})\to 0$. On the other hand, using the triangle inequality yields ${\rm RSE}(\mu)=W_{2}(\mu,\mathcal{E})\leq W_{2}(\nu_{i},\mu)+W_{2}(\nu_{i},\mathcal{E})$. Letting $i$ tend to infinity, we deduce ${\rm RSE}(\mu)\leq r$, which is a contradiction. Thus no such sequence $\nu_{i}$ exists and (12) holds.
Next, suppose that for some $c,q>0$, the inequality ${\rm RSE}(\nu)^{q}\leq c\cdot{\rm REG}(\nu)^{-1}$ holds for all $\nu\in\mathcal{Q}^{\prime}\setminus\mathcal{E}$. Then for any $\nu\in\mathcal{Q}^{\prime}$ such that $W_{2}(\nu,\mu)\leq r$, the triangle inequality yields
@@
Rearranging yields the estimate ${\rm REG}(\nu)\leq c\cdot({\rm RSE}(\mu)-r)^{-q}$, thereby completing the proof.